OpenStudy (anonymous):

solve for x cos(5x)+cos(2x)=0 all step answer Pi/7 + 4pi/7 k, 3pi/7+ 4pi/7 k, pi/3 + 4pi/3 k, where k=1,2,3,4,....

OpenStudy (experimentx):

cos 5x = -cos 3x = cos (pi + 3x)

OpenStudy (foolaroundmath):

Hint: Use \[\cos{C} + \cos{D} = 2\cos{\left(\frac{C+D}{2} \right)}\sin{\left(\frac{C-D}{2}\right)}\]

OpenStudy (anonymous):

Can you show all the steps?

OpenStudy (anonymous):